OpenStudy (brittany):

Write the equation in slope-intercept form of the line that contains the points (4, -7) and (0, 5).

OpenStudy (anonymous):

slope is -3

OpenStudy (anonymous):

then use y - y1 = m (x - x1)

OpenStudy (anonymous):

Slope: 5--7/0-4 12/-4 = -3 y = -3x + b 5 = 0 + b b = 5 y = -3x + 5

OpenStudy (curry):

y=-3x+5

OpenStudy (anonymous):

so y - 5 = -3 ( x - 0) y -5 = -3x y = -3x +5